OpenStudy (anonymous):

Which of the following is an equivalent form of the mixed number -12 3/4? a)45/4 b)-51/4 c)-45/4 d)-48/4

OpenStudy (ivettef365):

multiply (4)(-12) and add 3 and that result over 4

OpenStudy (anonymous):

\[-\left(12+\frac{3}{4}\right)=-\frac{51}{4} \]

OpenStudy (anonymous):

oh wait rob is right if you have to distribute the - nvrmnd

OpenStudy (anonymous):

kay, thanks everyone!